September 7, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, Septemeber 7, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

  • Updated error matrices
    • BCAL (Mark D.)
    • FCAL (Adesh/Matt)
    • Tracking (Simon)
  • FCAL timing updates (Adesh)
  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • FDC per-channel calibrations (Alex A./Mike S./?)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)
  • TOF alignment (Beni/Simon)


  1. Announcements
  2. Collaboration Meeting
  3. Run Planning
  4. Calibration Updates
  5. Simulations
    • Data/Simulation comparisons
      • BCAL (Will M.)
      • FCAL (Adesh)
      • CDC (Mike S.)
      • FDC (Alex A.)
      • Tracking (Paul/Simon)
      • TOF (Beni)
      • SC (Mahmoud)
  6. AOB


Attending: Sean (NU); Simon, Adesh, Mark I., Elton, Luke, Alex A., Paul, Beni, Nathan, Eugene (JLab); Naomi, Mike S., WIll M. (CMU); Matt S. (IU)

  1. Announcements
    • Alex A. reported that the Second Production Run is running full speed ahead, with 65% of batch 1 completed, which is expected to be completed by Friday. Production started Saturday afternoon after some last minute enhancements and bug fixes, and has run with up to ~6000 cores utilized.
    • Elton reported that the solenoid was ramped up to 1350A and was running fine as of 10 minutes before the meeting. If all goes well, this will be the planned setpoint for the Fall run.
    • Sean reported that the new EVIO parser had shown some minor potential issues with the EVIO skimming code, which he is tracking down before committing his changes. As reported previously, these changes enable targeted skimming of the data which drastically reduce skim size on disk and increase performance.
  2. Collaboration Meeting
    • Sean will give a calibration overview talk, and a requested talk on detector performance will be planned as well. The schedule will be firmed up at the offline software meeting.
  3. Run Planning
    • Data taking
      • Cosmic data taking is waiting on Alex S.'s upgrading of the firmware on all crates. The latest update is that it should be done by this coming Monday.
      • About 2 hours of Beam data at the start of the fall run is planned to be taken to recommission calibrations and the new data formats.
    • The output of the new fADC250 firmware will need to be tested offline. The latest news is that David has a spec to implement in the EVIO parser(s).
    • ROOT 6 transition - no one has tested their code with ROOT 6 yet. Mark I. said that there may be problems with his JLab builds, and is investigating. Sean had noticed that RooFIt was not being installed, although it was apparently enabled in the build.
    • Sean put together a list of currently existing calibrations for the fall run. We will review the readiness of these procedures at the next calibration meeting. [n.b. Sean had to leave the meeting due to a fire drill around this point.]
  4. Calibration Updates
    • Adesh is continuing to work on the FCAL timing calibration -- this is still a work in progress.
    • Andrei is working on the BCAL timing and the calibration for the effective velocity. He presented his procedure during a special calorimeter meeting; there is some documentation here.
    • Nathan is working on new iterative procedure for hodoscope timing
    • We had some discussion about how frequently calibrations for each of the detector systems changed.
      • Naomi pointed out that the CDC calibrations depend on temperature/pressure variations and so need to be monitored carefully and possibly modified with some frequency.
      • Lubomir indicated that the FDC calibrations do not need to changed frequently at all.
      • Adesh indicated that whenever the HV for the FCAL is updated a new calibration is needed; this procedure requires pi0 skims so cannot really be done "online"
    • Simon described recent changes to the tracking reconstruction code to improve track finding in the forward direction and changes to how the CDC and FDC geometries are handled to address strange phi-dependent distributions in the pull distributions resulting from kinematic fits. There still may be some residual effects.
    • Lubomir has been studying the behavior of the CDC for the transition region between the CDC and FDC at 10-15 degrees where hits on track are shared between the FDC and the CDC using data. He reports the CDC hit resolution seems to be poorer than expected for tracks in this angular range.
  5. Simulations
    • Data/Simulation comparisons
      • Nothing was reported for this yet.
  6. AOB